### v1

@@ 4.1 Overview
The general principle underlying RG-PT is that the reliability of some hyperparameters can be highly predictive of the reliability of other hyperparameters. This structure is encoded in an RG by leveraging prior information and data via the BT model Hunter (2004). By testing on a DAG, RG-PT supports a more flexible and efficient testing procedure that allows for the testing of multiple hyperparameters in parallel.
Using a partition $\{\mathcal{Z}_{\text{OPT}},\mathcal{Z}_{\text{MHT}}\}$ of the data set $\mathcal{Z}$, at a high level, RG-PT consists of the following steps, which are illustrated in Fig. 4:
Estimating the Pareto front for all risk measures: Using the data set $\mathcal{Z}_{\text{OPT}}$, RG-PT identifies the the subset $\Lambda_{\text{OPT}}\subseteq\Lambda$ of hyperparameters that are estimated to lie on the Pareto front in the same manner as PT (see previous section).
Learning the reliability graph: Rather than ordering the hyperparameters globally and sequentially as done by PT, RG-PT creates an RG, with nodes given by the hyperparameters in subset $\Lambda_{\text{OPT}}$. This is done by following the principle that hyperparameters $\lambda\in\Lambda_{\text{OPT}}$ whose reliability levels are predictive of the reliability levels of other hyperparameters $\Lambda^{\prime}\subset\Lambda_{\text{OPT}}$ should be tested before the hyperparameters $\Lambda^{\prime}$. Details are provided in Sec. 4.2.
FDR-controlling MHT: FDR-controlling MHT is carried out by following the order encoded by the RG. As explained in Sec. 4.2, this is done by using DAGGER Ramdas et al. (2019), targeting FDR requirements.
Addressing the multi-objective optimization problem: Using the estimated subset $\hat{\Lambda}_{\mathcal{Z}}$ of reliable hyperparameters, the multi-objective optimization problem (7) is addressed in the same manner as PT (see previous section).
Figure 4: Illustration of the main steps of RG-PT: ① Estimate the hyperparameters $\Lambda_{\text{OPT}}$ lying on the Pareto front; ② Build the RG over the selected hyperparameters; ③ Apply an FDR-controlling MHT procedure, DAGGER, to the RG.
@@ 4.2 Learning the Reliability Graph
After obtaining the estimated Pareto front $\Lambda_{\text{OPT}}$, RG-PT constructs an RG to encode the expected relationships between the reliability levels attained by the candidate hyperparameters in the set $\Lambda_{\text{OPT}}$. The RG is a DAG, in which each node represents a hyperparameter, and edges are directed to describe a reliability hierarchy. Specifically, edges encode the expectation that parent nodes are predictive of the reliability of their child nodes.
Starting from the nodes with no parents and following the direction of the edges in the RG, one encounters hyperparameters that are estimated to be increasingly unreliable. Generalizing the linear ordering assumed by PT, the DAG structure adopted by RG-PT can assign the same expected reliability ranking to multiple hyperparameters. Specifically, all the hyperparameters at the same depth in the DAG are deemed to have the same relative reliability level.
In order to learn the RG, RG-PT leverages the data set $\mathcal{Z}_{\text{OPT}}$, as well as, possibly, prior information about the relative reliability of pairs of hyperparameters. This is done by following two steps:
1. Depth assignment: The hyperparameters in set $\Lambda_{\text{OPT}}$ are ranked in terms of their expected reliability, while allowing for multiple hyperparameters to be ranked equally. This effectively assigns a depth level $d$ in the DAG to each hyperparameter, with multiple hyperparameters possibly sharing the same depth level. This step is explained in Sec. 4.2.1.
2. Learning the directed edges: Given any hyperparameter $\lambda$ at some depth $d$, RG-PT selects a subset of hyperparameters at the previous depth level $d-1$ to serve as parents of the hyperparameter $\lambda$. This is done by choosing the hyperparameters at depth $d-1$ that are most predictive of the reliability level of hyperparameter $\lambda$. This step is explained in Sec. 4.2.2.
@@ 4.2.1 Depth Assignment
To compute the scores $s(\lambda)$ for hyperparameters $\lambda\in\Lambda_{\text{OPT}}$, we use the BT model Hunter (2004). The BT model converts pairwise counts $w_{ij}$ for all pairs of hyperparameters $\lambda_{i}$ and $\lambda_{j}$ in subset $\Lambda_{\text{OPT}}$ into per-hyperparameter scores $s(\lambda)$ for all $\lambda\in\Lambda_{\text{OPT}}$. The pairwise counts measure the number of times that hyperparameter $\lambda_{i}$ was found to be more reliable than hyperparameter $\lambda_{j}$. In RG-PT, we propose to evaluate the pairwise counts $w_{ij}$ by leveraging two sources of information:
$\bullet$ Prior information: Prior information is encoded by pairwise probabilities $0\leq\eta_{ij}\leq 1$ for each pair of hyperparameters $\lambda_{i},\lambda_{j}\in\Lambda_{\text{OPT}}$. This probability reflects the expected rate at which hyperparameter $\lambda_{i}$ is observed to be more reliable than hyperparameter $\lambda_{j}$. Note that we have $\eta_{ji}=1-\eta_{ij}$. The strength of the prior information is determined via a pseudocount variable $n_{p}$ as in the standard categorical-Dirichlet model Bishop & Nasrabadi (2006). A larger pseudocount $n_{p}$ indicates a stronger trust in the prior information. Importantly, the statistical guarantees of RG-PT do not depend on the choice of the prior probabilities $\{\eta_{ij}\}$ and pseudocount $n_{p}$, which can, however, improve the capacity of RG-PT to optimize the auxiliary risk functions. Note that in the absence of prior information, one can set $n_{p}=0$.
@@
Overall, the pair-wise probability $w_{ij}$ is obtained by combining prior information and data as
@@
so that the strength of the prior depends on the ratio $n_{p}/|\mathcal{Z}_{\text{OPT}}|$ between the pseudocount $n_{p}$ and the number of data points $|\mathcal{Z}_{\text{OPT}}|$.
Using the BT model, the scores $s(\lambda_{i})$ for all hyperparameters $\lambda_{i}\in\Lambda_{\text{OPT}}$ are obtained by maximizing the log-likelihood Hunter (2004)
@@
with $w_{ii}=0$ for all $1\leq i\leq|\Lambda_{\text{OPT}}|$. With this design, in the absence of prior information ($n_{p}=0$), the BT model reduces to assigning scores directly proportional to the $p$-values $p_{\lambda}(\mathcal{Z}_{\text{OPT}})$, as is done in PT Laufer-Goldshtein et al. (2023). Therefore, the proposed approach ensures a seamless integration of prior information into the score assignment, while maintaining consistency with PT when prior information is unavailable.
@@
Clustering can be implemented by using methods such as $K$-means or hierarchical clustering. We recommend to use agglomerative hierarchical clustering, which begins with each hyperparameter in its own cluster and iteratively merges clusters Jain & Dubes (1988).
@@ 4.2.2 Learning the Directed Edges
To this end, we implement feature selection via the non-negative Lasso Tibshirani (1996). Specifically, given hyperparameter $\lambda\in\Lambda_{d}$, we consider the problem of predicting the risks $\{r_{l}(Z,\lambda)\}_{l=1}^{L}$ from the risks $\{r_{l}(Z,\lambda^{\prime})\}_{l=1}^{L}$ attained by the hyperparameters $\lambda\in\Lambda_{d-1}$ at the previous depth level. The use of non-negative Lasso regression ensures that only positive correlations are represented in the DAG, preserving hierarchical reliability relationships between parent and child nodes.
@@
$\min_{\beta\geq 0}$ $\sum_{Z\in\mathcal{Z}_{\text{OPT}}}\|r(Z,\lambda)-\sum_{\lambda^{\prime}\in\Lambda_{d-1}}\beta_{\lambda^{\prime}}r(Z,\lambda^{\prime})\|_{2}^{2}$ (13)
$+\tau\sum_{\lambda^{\prime}\in\Lambda_{d-1}}\beta_{\lambda^{\prime}},$
where $\beta=\{\beta_{\lambda^{\prime}}\}_{\lambda^{\prime}\in\Lambda_{d-1}}$ is the vector of non-negative regression coefficients corresponding to each potential parent node $\lambda^{\prime}\in\Lambda_{d-1}$; $r(Z,\lambda)$ is the vector containing the values $\{r_{l}(Z,\lambda)\}_{l=1}^{L_{c}}$; $\|\cdot\|_{2}$ represents the $\ell_{2}$ norm; and $\tau>0$ is a regularization parameter that controls the degree of sparsity in the solution. As for the variables $(n_{p},\{\eta_{ij}\})$ in the BT likelihood (12), the choice of the parameter $\tau$ does not affect the validity properties of RG-PT. After solving the convex problem (13), only the hyperparameters $\lambda^{\prime}\in\Lambda_{j-1}$ for which the corresponding coefficient $\beta_{\lambda^{\prime}}$ are positive are selected as parent nodes of hyperparameter $\lambda$.
@@ 4.3 FDR-Controlling Multiple Hypothesis Testing
Given the obtained RG, RG-PT performs MHT via DAGGER Ramdas et al. (2019), an FDR-controlling algorithm that operates on DAGs. Next, we first outline the reasons behind using FDR, and then explain the testing procedure.
The requirement (6) guaranteed by FWER-controlling algorithms often comes at the cost of limiting the number of hyperparameters retained in the subset $\hat{\Lambda}_{\mathcal{Z}}$. This, in turn, constrains the quality of the solutions to problem (7). Instead of constraining the probability of having any unreliable hyperparameter in set $\hat{\Lambda}_{\mathcal{Z}}$, one can control the proportion of unreliable hyperparameters in set $\hat{\Lambda}_{\mathcal{Z}}$.
Accordingly, the FDR is defined as the expected proportion of unreliable hyperparameters in set $\hat{\Lambda}_{\mathcal{Z}}$. Controlling the FDR amounts to finding a subset $\hat{\Lambda}_{\mathcal{Z}}\subseteq\Lambda$ that satisfies the inequality
where $\mathbf{1}\{\cdot\}$ is the indicator function, and the expectation is taken over the unknown data distribution $P_{Z}$. The FDR constraint (14) contrasts with the FWER requirement (6), which aims to limit the probability of any unreliable hyperparameter in the selected set $\hat{\Lambda}_{\mathcal{Z}}$.
DAGGER is an FDR-controlling procedure that operates on a DAG. DAGGER begins testing at the root nodes of the RG, i.e., at the hyperparameters in cluster $\Lambda_{1}$, and proceeds with the clusters $\Lambda_{2},\Lambda_{3},\ldots,\Lambda_{D}$, guided by the outcomes of prior tests. If a hyperparameter is deemed unreliable, none of its descendants are tested.
Using the data set $\mathcal{Z}_{\text{MHT}}$, the $p$-values used for MHT are given by $p_{\lambda}(\mathcal{Z}_{\text{MHT}})$ in (9) for each hyperparameter $\lambda\in\Lambda_{\text{OPT}}$. The testing level $\delta_{i}$ for each hyperparameter $\lambda_{i}$, detecting $\lambda_{i}$ as reliable if $p_{\lambda_{i}}\leq\delta_{i}$, is determined based on several factors. These include the overall target FDR level $\delta$ in constraint (14), the number of reliable hyperparameters identified among those tested prior to $\lambda_{i}$, and the structure of the graph rooted at $\lambda_{i}$. We refer the reader to Appendix B and to Ramdas et al. (2019) for details.
@@ 5 Experiments
In this section, we evaluate the proposed RG-PT hyperparameter selection strategy on two problems: a sequence-to-sequence translation problem Peters & Martins (2021), and an object detection task Angelopoulos et al. (2021). Two additional experiments addressing a classical small-scale machine learning benchmark Franceschi et al. (2024) and an engineering application in the domain of telecommunications Valcarce (2020) can be found in Appendix C [^1] The codes for the experiments can be found at https://github.com/kclip/RG-PT..
Throughout, we adopt as benchmarks LTT Angelopoulos et al. (2021) and PT Laufer-Goldshtein et al. (2023). LTT is implemented by applying Benjamini-Hochberg (BH) Benjamini & Hochberg (1995) as the FDR-controlling algorithm, while PT follows Sec. 3 with the caveat that FDR-controlling FST Lynch et al. (2017) is used in lieu of an FWER-controlling scheme. LTT uses the entire calibration data set $\mathcal{Z}$ to evaluate the $p$-values used in BH, while PT and RG-PT partition $\mathcal{Z}$ into data sets $\mathcal{Z}_{\text{OPT}}$ and $\mathcal{Z}_{\text{MHT}}$. To the best of our knowledge, LTT and PT are the only existing hyperparameter selection methods that guarantee statistical validity in the sense of the FWER requirement (6) or the FDR constraint (14).

### v2

@@ 3.1 Overview
The design of RG-PT starts from the observation that the reliability of some hyperparameters can be highly predictive of the reliability of other hyperparameters, and that this structure can be encoded by a DAG as in Fig. 1(a). By incorporating the DAG structure in the MHT process of hyperparameter selection, RG-PT supports a more efficient hyperparameter selection procedure, while meeting formal reliability constraints in terms of the FDR (8).
As illustrated in Fig. 2, using a partition $\mathcal{Z}=\{\mathcal{Z}_{\text{OPT}},\mathcal{Z}_{\text{MHT}}\}$ of the data set $\mathcal{Z}$, RG-PT applies the following steps:
Estimating the Pareto front for all risk measures: Following PT [5], RG-PT uses the dataset $\mathcal{Z}_{\text{OPT}}$ to identify the subset $\Lambda_{\text{OPT}}\subseteq\Lambda$ of hyperparameters that are on the Pareto front of the space of estimated risk measures $\{\hat{R}_{l}(\lambda|\mathcal{Z}_{\text{OPT}})\}_{l=1}^{L}$. This is done by addressing the multi-objective optimization problem (3) with the estimates $\{\hat{R}_{l}(\lambda|\mathcal{Z}_{\text{OPT}})\}_{l=1}^{L}$ in lieu of the true risks $\{R_{l}(\lambda|\mathcal{Z}_{\text{OPT}})\}_{l=1}^{L}$ using any suitable multi-objective optimization algorithm [5].
Figure 2: Illustration of the main steps of RG-PT: ① Estimate the hyperparameters $\Lambda_{\text{OPT}}$ lying on the Pareto front pf problem (2); ② Build the RG over the selected hyperparameters $\Lambda_{\text{OPT}}$; ③ Apply an FDR-controlling MHT procedure, DAGGER, to the RG to obtain the selected set $\hat{\Lambda}_{\mathcal{Z}}\subseteq\Lambda_{\text{OPT}}$.
Learning the reliability graph: Rather than ordering the hyperparameters in subset $\Lambda_{\text{OPT}}$ in a linear sequence as done by PT [5], RG-PT creates an RG, with nodes given by the hyperparameters in subset $\Lambda_{\text{OPT}}$. This is done by following the principle that hyperparameters $\lambda\in\Lambda_{\text{OPT}}$ whose reliability levels are predictive of the reliability levels of other hyperparameters $\Lambda^{\prime}\subset\Lambda_{\text{OPT}}$ should be tested before the hyperparameters $\Lambda^{\prime}$. As detailed in Sec. 3.2, the RG construction leverages the BT model to incorporate prior information and the non-negative Lasso to determine the links in the graph.
FDR-controlling MHT: Using the data set $\mathcal{Z}_{\text{MHT}}$, FDR-controlling MHT is carried out by incorporating the structure encoded by the RG. As explained in Sec. 3.2, this is done by using DAGGER [15], returning the subset $\hat{\Lambda}_{\mathcal{Z}}\subseteq\Lambda_{\text{OPT}}$.
Addressing the multi-objective optimization problem: Given the subset $\hat{\Lambda}_{\mathcal{Z}}$, RG-PT addresses the problem
@@ 3.2 Learning the Reliability Graph
After obtaining the estimated Pareto front $\Lambda_{\text{OPT}}$, RG-PT constructs an RG to encode the expected relationships between the reliability levels attained by the candidate hyperparameters in the set $\Lambda_{\text{OPT}}$. The RG is a DAG in which each node represents a hyperparameter $\lambda\in\Lambda_{\text{OPT}}$, and edges are directed to describe a reliability hierarchy. Specifically, edges encode the expectation that parent nodes are predictive of the reliability of their child nodes.
Accordingly, starting from the nodes with no parents and following the direction of the edges in the RG, one encounters hyperparameters that are estimated to be increasingly unreliable. Generalizing the linear ordering assumed by PT, the DAG structure adopted by RG-PT can thus assign the same expected reliability ranking to multiple hyperparameters. Specifically, all the hyperparameters at the same depth in the DAG are deemed to have the same relative reliability level. This partial ordering enables a more efficient exploration of the hyperparameter space.
In order to learn the RG, RG-PT leverages the data set $\mathcal{Z}_{\text{OPT}}$, as well as, possibly, prior information about the relative reliability of pairs of hyperparameters. This is done via the following two steps:
1. Depth assignment: The hyperparameters in set $\Lambda_{\text{OPT}}$ are ranked in terms of their expected reliability, allowing for multiple hyperparameters to be ranked equally. This step thus assigns a depth level $d$ in the DAG to each hyperparameter $\lambda\in\Lambda_{\text{OPT}}$. As explained in Sec. 3.2.1, this is done by leveraging the BT ranking model [13].
2. Learning the directed edges: Given any hyperparameter $\lambda$ at some depth $d$, RG-PT selects a subset of hyperparameters at the previous depth level $d-1$ to serve as parents of the hyperparameter $\lambda$. As specified in Sec. 3.2.2, this is done by choosing the hyperparameters at depth $d-1$ that are most predictive of the reliability level of hyperparameter $\lambda$ via the non-negative Lasso [14].
@@ 3.2.1 Depth Assignment
To compute the scores $s(\lambda)$ for hyperparameters $\lambda\in\Lambda_{\text{OPT}}$, we use the BT model [13]. The BT model converts pairwise counts $w_{ij}$ for all pairs of hyperparameters $\lambda_{i}$ and $\lambda_{j}$ in subset $\Lambda_{\text{OPT}}$ into per-hyperparameter scores $s(\lambda)$ for all $\lambda\in\Lambda_{\text{OPT}}$. The pairwise count $w_{ij}$ measures the number of times that hyperparameter $\lambda_{i}$ was found to be more reliable than hyperparameter $\lambda_{j}$. In RG-PT, we propose to evaluate the pairwise counts $w_{ij}$ by leveraging two sources of information:
$\bullet$ Prior information: Prior information is encoded by pairwise probabilities $0\leq\eta_{ij}\leq 1$ for each pair of hyperparameters $(\lambda_{i},\lambda_{j})$ in $\Lambda_{\text{OPT}}$. This probability reflects the expected rate at which hyperparameter $\lambda_{i}$ is observed to be more reliable than hyperparameter $\lambda_{j}$. Note that we have $\eta_{ji}=1-\eta_{ij}$. The strength of the prior information is determined via a pseudocount variable $n_{p}$ as in the standard categorical-Dirichlet model [25]. A larger pseudocount $n_{p}$ indicates a stronger trust in the prior information. Importantly, the statistical guarantees of RG-PT do not depend on the choice of the prior probabilities $\{\eta_{ij}\}$ and pseudocount $n_{p}$, which can, however, improve the capacity of RG-PT to optimize the auxiliary risk functions. Note that in the absence of prior information, one can set $n_{p}=0$.
@@
Overall, the pair-wise count $w_{ij}$ is obtained by combining prior information and data as
@@
so that the relative strength of the prior information in (11) depends on the ratio $n_{p}/|\mathcal{Z}_{\text{OPT}}|$ between the pseudocount $n_{p}$ and the number of data points $|\mathcal{Z}_{\text{OPT}}|$.
Using the BT model, the scores $s(\lambda_{i})$ for all hyperparameters $\lambda_{i}\in\Lambda_{\text{OPT}}$ are obtained by maximizing the log-likelihood [13]
@@
with $w_{ii}=0$ for all $1\leq i\leq|\Lambda_{\text{OPT}}|$. With this design, in the absence of prior information ($n_{p}=0$), the BT model reduces to assigning scores directly proportional to the $p$-values $p_{\lambda}(\mathcal{Z}_{\text{OPT}})$.
@@
Clustering can be implemented by using methods such as $K$-means or hierarchical clustering. We recommend using agglomerative hierarchical clustering, which begins with each hyperparameter in its own cluster and iteratively merges clusters [26].
@@ 3.2.2 Learning the Directed Edges
To this end, we implement feature selection via the non-negative Lasso [14]. Specifically, given hyperparameter $\lambda\in\Lambda_{d}$, we consider the problem of predicting the risks $\{r_{l}(Z,\lambda)\}_{l=1}^{L}$ from the risks $\{r_{l}(Z,\lambda^{\prime})\}_{l=1}^{L}$ attained by the hyperparameters $\lambda\in\Lambda_{d-1}$ at the previous depth level. The use of non-negative Lasso regression ensures that only positive correlations are represented in the DAG, preserving hierarchical reliability relationships between parent and child nodes.
@@
$\min_{\beta\geq 0}\sum_{Z\in\mathcal{Z}_{\text{OPT}}}\left\|r(Z,\lambda)-\sum_{\lambda^{\prime}\in\Lambda_{d-1}}\beta_{\lambda^{\prime}}r(Z,\lambda^{\prime})\right\|_{2}^{2}+\tau\sum_{\lambda^{\prime}\in\Lambda_{d-1}}\beta_{\lambda^{\prime}},$ (13)
where $\beta=\{\beta_{\lambda^{\prime}}\}_{\lambda^{\prime}\in\Lambda_{d-1}}$ is the vector of non-negative regression coefficients corresponding to each potential parent node $\lambda^{\prime}\in\Lambda_{d-1}$; $r(Z,\lambda)$ is the vector containing the values $\{r_{l}(Z,\lambda)\}_{l=1}^{L_{c}}$; $\|\cdot\|_{2}$ represents the $\ell_{2}$ norm; and $\tau>0$ is a regularization parameter that controls the degree of sparsity in the solution. After solving the convex problem (13), only the hyperparameters $\lambda^{\prime}\in\Lambda_{j-1}$ for which the corresponding coefficient $\beta_{\lambda^{\prime}}$ are positive are selected as parent nodes of hyperparameter $\lambda$. As for the variables $(n_{p},\{\eta_{ij}\})$ in the BT likelihood (12), the choice of the parameter $\tau$ does not affect the validity properties of RG-PT.
@@ 3.3 FDR-Controlling Multiple Hypothesis Testing
Given the obtained RG, RG-PT performs MHT via DAGGER [15], an FDR-controlling algorithm that operates on DAGs. DAGGER begins testing at the root nodes of the RG, i.e., at the hyperparameters in cluster $\Lambda_{1}$, and proceeds with the clusters $\Lambda_{2},\Lambda_{3},\ldots,\Lambda_{D}$, guided by the outcomes of prior tests. If a hyperparameter is deemed unreliable, none of its descendants are tested.
The test for each hyperparameter $\lambda_{i}\in\Lambda_{\text{OPT}}$ detects $\lambda_{i}$ as reliable if the p-value $p_{\lambda_{i}}(\mathcal{Z}_{\text{MHT}})$ in (7) is no larger than a threshold $\delta_{i}$, i.e., if $p_{\lambda_{i}}(\mathcal{Z}_{\text{MHT}})\leq\delta_{i}$. The testing level $\delta_{i}$ is determined by DAGGER based on several factors, including the overall target FDR level $\delta$ in constraint (8), the number of reliable hyperparameters identified among those tested prior to $\lambda_{i}$, and the structure of the graph rooted at $\lambda_{i}$. We refer the reader to Appendix B and to [15] for details.
An algorithmic overview of RG-PT is provided in Appendix D. The following proposition states the theoretical guarantees provided by RG-PT.
Proposition 3.1.
The set $\hat{\Lambda}_{\mathcal{Z}}$ of hyperparameters returned by RG-PT controls the FDR below the pre-specified threshold $\delta$ as in (8).
Proof.
RG-PT applies the DAGGER algorithm [15] to a DAG over the candidate hyperparameters, using valid p-values defined in (7). Since DAGGER controls the FDR at level $\delta$ for any DAG structure when supplied with valid p-values under arbitrary dependence (see Appendix B), the result follows directly. ∎
@@ 4 Experiments
In this section, we evaluate the proposed RG-PT hyperparameter selection strategy on a prompt engineering problem [1] and a sequence-to-sequence translation task [27]. Additional experiments including on object detection [4] and telecommunications [28] can be found in Appendix E [^1] The code for the experiments can be found at the Github repository https://github.com/kclip/RG-PT.
Throughout the experiments, we adopt as benchmarks LTT [4] and PT [5]. To the best of our knowledge, LTT and PT are the only existing hyperparameter selection methods that guarantee statistical validity in the sense of the FDR constraint (8), justifying this choice. LTT is implemented by applying Benjamini-Hochberg (BH) [29] as the FDR-controlling algorithm, while PT follows [5], with the caveat that FDR-controlling FST [30] is used in lieu of an FWER-controlling scheme. LTT uses the entire calibration data set $\mathcal{Z}$ to evaluate the p-values used in BH, while PT and RG-PT partition $\mathcal{Z}$ into data sets $\mathcal{Z}_{\text{OPT}}$ and $\mathcal{Z}_{\text{MHT}}$.
4.1 Reliable Prompt Engineering
Problem Setup. In this experiment, we focus on prompt engineering for the following three tasks from the instruction induction data set [1]: 1. Sentiment analysis: In this task, based on the Stanford Sentiment Treebank dataset [6], each data point $Z=(X,Y)$ encompasses a movie review $X$, and the corresponding sentiment $Y\in\{\text{positive},\text{negative}\}$. 2. Sentence similarity: In this task, based on the Semantic Textual Similarity Benchmark dataset [31], each data point $Z=(X,Y)$ comprises two sentences as input $X$, along with a semantic similarity label $Y\in\{0,\dots,5\}$. 3. Word in context: In this task, based on the Word-in-Context dataset [32], each data point $Z=(X,Y)$ consists of a target word and two context sentences as input $X$, paired with a binary label $Y\in\{\text{same},\text{not same}\}$ indicating whether the target word shares the same meaning across both contexts.
For each task, we use 1000 examples each for the data sets $\mathcal{Z}_{\text{OPT}}$ and $\mathcal{Z}_{\text{MHT}}$, as well as for the test data set. Furthermore, following the forward generation mode detailed in [33], we use the LLaMA3-70B-Instruct model [34] to generate a set $\Lambda=\{\lambda_{1},\dots,\lambda_{100}\}$ of distinct instruction-style prompt templates for each task.
Given a prompt $\lambda$ and an input $X$, the smaller LLaMA3-8B-Instruct model [35] $f$ is queried with the concatenated input $[\lambda,X]$, producing the output $f([\lambda,X])$. For each input-output pair $Z=(X,Y)$, a task-specific 0-1 prompt loss is calculated as $r_{\text{prompt}}(Z,\lambda)=l(f([\lambda,X]),Y)\in\{0,1\}$, indicating whether the task was performed correctly. The objective is to find prompts in set $\Lambda$ that control the average prompt loss $R_{\text{prompt}}(\lambda)=\mathbb{E}_{\mathcal{Z}}\left[r_{\text{prompt}}(Z,\lambda)\right]$ below a target level of $\alpha=0.2$, while minimizing the average prompt length. For this selection, we wish to control the FDR in (8) at level $\delta=0.1$.
[…]
Results. For each task, we plot the distribution of the length of the shortest reliable prompt for 100 independent runs over random splits of the data set, for LTT, PT, and RG-PT in Fig. 1(b), Fig. 3(a), and Fig. 3(b) for the sentiment analysis task, the sentence similarity task, and the word in context task, respectively. The figures demonstrate that RG-PT identifies more concise instructions compared to LTT and PT by leveraging the prior information provided by the LLM judge. Note that all schemes satisfy the FDR constraint (not shown). For instance, RG-PT achieved an average FDR of 0.089, 0.095, and 0.092 for the the sentiment analysis, the sentence similarity, and the word in context tasks, respectively.
Figure 3: Distribution of the length of the shortest prompt templates identified by LTT [4], PT [5], and the proposed RG-PT for (a) the Semantic Textual Similarity Benchmark dataset [31] and (b) the Word-in-Context dataset [32].
An ablation study on the effect of the RG depth $D$, as well as the effect of a misspecified prior information for this experiment can be found in Appendix E.3.